### v2

# Fremlin tensor product respects the unbounded order convergence

Abstract.
Suppose $\Sigma$ is a topological space and $S(\Sigma)$ is the vector lattice of all equivalence classes of continuous real-valued functions defined on open dense subsets of $\Sigma$. In this paper, we establish some lattice and topological aspects of $S(\Sigma)$. In particular, as an application, we show that the unbounded order convergence and the order convergence are stable under passing to the Fremlin tensor product of two Archimedean vector lattices. Furthermore, by considering the Fremlin projective tensor product between Banach lattices, we establish a partial variant of this result for unbounded norm convergence and unbounded absolute weak convergence, as well.
Key words and phrases:
Fremlin tensor product, unbounded order convergence, order convergence, vector lattice.
2020 Mathematics Subject Classification:
Primary: 46M05. Secondary: 46A40.
July 3, 2024
1. Introduction
Suppose $E$ and $F$ are Archimedean vector lattices. Fremlin in [9] has constructed the Archimedean vector lattice $E\overline{\otimes}F$ (known as the Fremlin tensor product of two Archimedean vector lattices). Under this structure, the algebraic tensor product $E\otimes F$ can be considered as an ordered vector subspace of $E\overline{\otimes}F$. There are various properties of $E$ and $F$ that can be or can not be inherited by the Fremlin tensor product. For example, the Fremlin tensor product behaves well with sublattices; that is if $E_{0}$ and $F_{0}$ are vector sublattices of vector lattices $E$ and $F$, respectively, then $E_{0}\overline{\otimes}F_{0}$ is a vector sublattice of $E\overline{\otimes}F$. Furthermore, if $E_{0}$ and $F_{0}$ are order dense or regular, then, so is the Fremlin tensor product $E_{0}\overline{\otimes}F_{0}$ (for more details, see[13, Proposition 3.2]). However, there are many properties that fail in the Fremlin tensor product between vector lattices even so the underlying spaces possess the corresponding properties, as well. For example, order completeness can not be preserved by the Fremlin tensor product. More precisely, Fremlin in [10, 4C], showed that the Fremlin tensor product of $L^{2}[0,1]$ with itself fails to be order complete even when completed relative to its norm. Also, Grobler in [13, Example 3.8], constructed an example of two compact Hausdorff extremally disconnected topological spaces $K_{1}$ and $K_{2}$ such that $K_{1}\times K_{2}$ is not extremally disconnected so that $C(K_{1}\times K_{2})$ is not order complete, even so both $C(K_{1})$ and $C(K_{2})$ are order complete.
Moreover, the Fremlin tensor product does not respect ideals in general (see [7, Theorem 3.8] for more details). In fact, the authors in [7] have constructed an ideal whose the Fremlin tensor product with itself is not an ideal. Since tensor products are a useful tool in studying other mathematical subjects, it is interesting in its own right, to see whether or not some properties of the underlying spaces can be inherited by the tensor product. One of the possible questions arises while we are dealing with different convergence notions. For example, it is known that the projective tensor norm on the projective tensor product of two Banach spaces is a cross-norm so that it preserves the norm convergence. However, it does not behave well with the weak convergence. More precisely, it is shown in [15, Example 2.10] that the diagonal of the projective tensor product $\ell_{2}\widehat{\otimes}\ell_{2}$ is isometrically isomorphic to $\ell_{1}$. Note that by diagonal, we mean the closed subspace of $\ell_{2}\widehat{\otimes}\ell_{2}$ generated by the elementary tensors $e_{n}\otimes e_{n}$, in which $(e_{n})_{n\in\mathbb{N}}$ is the standard basis of $\ell_{2}$. Moreover, if $u=\Sigma_{n=1}^{\infty}\alpha_{n}e_{n}\otimes e_{n}$, then it can be verified that the projective norm of $u$, denoted by $\pi(u)$, is equal to $\Sigma_{n=1}^{\infty}|\alpha_{n}|$ so that $D$ is isometrically isomorphism to $\ell_{1}$. Now consider the standard basis $(e_{n})\subseteq\ell_{2}$ which is weakly null but $e_{n}\otimes e_{n}$ is not since $\pi(e_{n}\otimes e_{n})=1$ and in $\ell_{1}$, weakly convergence and norm convergence agree. It is known that, in the case of a vector lattice, we have two convergence structures: order convergence and unbounded order convergence. Therefore, it would be interesting to consider the following question:
Question 1.
Does the Fremlin tensor product of two vector lattices preserve either order convergence or unbounded order convergence?
Grobler in [13, Corollary 3.4] considered a ”separate” version answer to this question. More precisely, he proved that for Archimedean vector lattices $E$ and $F$, if $\sigma:E\times F\to E\overline{\otimes}F$ is the natural lattice bimorphism ($(x,y)\rightarrow x\otimes y$), then $\sigma$ is order ($uo$-continuous) on each component. In Section 3, we show that the Fremlin tensor product of two Archimedean vector lattices behave well with both order convergence and unbounded order convergence in a general sense; this extends the result of Grobler to a ”jointly” version, as well. Before we reach to this goal, we need to verify some lattice and topological structures for the vector lattice of all equivalence classes of continuous real-valued functions defined on open dense subsets of a topological space $\Sigma$. Suppose now $E$ and $F$ are Banach lattices. Fremlin in [10] has constructed the Banach lattice $E\widehat{\otimes}F$ (known as the Fremlin projective tensor product of $E$ and $F$); $E\overline{\otimes}F$ is norm dense vector sublattice in $E\widehat{\otimes}F$ so that a normed lattice in its own right. On the other hand, while we are dealing with Banach lattices, we have two known natural notions for unbounded convergences: unbounded norm convergence and unbounded absolute weak convergence. Therefore, the following question is natural.
Question 2.
Does the Fremlin projective tensor product of two Banach lattices preserve either unbounded norm convergence or unbounded absolute weak convergence?
In Section 4, we shall provide a partial positive answer to this question, as well. In Section 2, we provide some preliminaries that are necessary for the main results. In fact, we give a brief exposition on unbounded convergences and also on the Fremlin tensor products, as well.
2. preliminaries
2.1. unbounded convergences
[…]
2.2. the Fremlin tensor products
In this part, we recall some notes about the Fremlin tensor product between vector lattices and also Banach lattices. For more details, see [9, 10]. Furthermore, for a comprehensive, new and interesting reference, see [18]. Furthermore, for a short and nicely written exposition on different types of tensor products between Archimedean vector lattices, see [13].
Suppose $E$ and $F$ are Archimedean vector lattices. In 1972, Fremlin constructed a tensor product $E\overline{\otimes}F$ that is an Archimedean vector lattice with the following properties:
  • The algebraic tensor product $E\otimes F$ is a vector subspace of $E\overline{\otimes}F$ so that it is an ordered vector subspace in its own right.
  • The vector sublattice in $E\overline{\otimes}F$ generated by $E\otimes F$ is the whole of $E\overline{\otimes}F$.
  • For each Archimedean vector lattice $G$ and every lattice bimorphism $\Phi:E\times F\to G$, there is a unique lattice homomorphism $T:E\overline{\otimes}F\to G$ such that $T(x\otimes y)=\Phi(x,y)$ for each $x\in E$ and for each $y\in F$.
Therefore, every element of $E\overline{\otimes}F$ can be considered as a finite suprema and finite infima of some elements of $E\otimes F$. The good news is that we have some density properties for elements of $E\overline{\otimes}F$ in terms of the elements of the algebraic tensor product $E\otimes F$ as follows.
  • Assume that $E$ and $F$ are Archimedean vector lattice. For each $a\in E\overline{\otimes}F$, there exist $u\in E_{+}$ and $v\in F_{+}$ such that for each $\varepsilon>0$, there is $b\in E\otimes F$ with $|a-b|\leq\varepsilon u\otimes v$, [18, Proposition 3.11].
  • For each $c\in(E\overline{\otimes}F)_{+}$, we have $c=\sup\{a\otimes b:a\in E_{+},b\in F_{+}\}$, [18, Proposition 3.12].
  • For each $u\in(E\overline{\otimes}F)_{+}$, there exist $a\in E_{+}$ and $b\in F_{+}$ with $u\leq a\otimes b$, [10, 1A(d)].
Now, suppose $E$ and $F$ are Banach lattices. Fremlin in [10] constructed a tensor product $E\widehat{\otimes}F$ that is a Banach lattice with the following properties.
  • $E\widehat{\otimes}F$ is the norm completion of $E\otimes F$ with respect to the positive-projective norm: for each $u=\Sigma_{i=1}^{n}x_{i}\otimes y_{i}\in E\otimes F$:
[…]
  • $E\overline{\otimes}F$ can be considered as an norm-dense vector sublattice of $E\widehat{\otimes}F$.
  • For every Banach lattice $G$ and every continuous positive bilinear mapping $\Phi:E\times F\to G$, there is a unique positive increasing linear mapping $T:E\widehat{\otimes}F\to G$ with $T(x\otimes y)=\Phi(x,y)$; moreover, this correspodence is norm-preserving.
  • The positive cone in $E\widehat{\otimes}F$ is the closure of the cone $P\subseteq E\otimes F$ generated by $\{x\otimes y:x\in E_{+},y\in F_{+}\}$.
  • The projective norm, .$\|.\|_{\pi}$, on $E\widehat{\otimes}F$ is a cross-norm that is for every $x\in E$ and for every $y\in F$, we have $\|x\otimes y\|_{\pi}=\|x\|\|y\|$.
3. main results
3.1. Fremlin tensor product and $uo$-convergence
Suppose $X$ is a topological space and consider the space $C^{\infty}(X)$ consisting of all extended continuous functions from $X$ into the extended reals ($\mathbb{R}^{*}=[-\infty,\infty]$). In general, this space need not be a vector space (see [6, Section 2]). However, when $X$ is extremally disconnected, it is known that $C^{\infty}(X)$ is a vector lattice under the pointwise vector and order operations. Suppose $K_{1}$ and $K_{2}$ are compact Hausdorff spaces. It is known that $C(K_{1})\otimes C(K_{2})$ is norm (order) dense in $C(K_{1}\times K_{2})$. On the other hand, by the known Kakutani’s theorem ([2, Theorem 4.21]), every Archimedean vector lattice with an order unit can be considered as a norm (order) vector sublattice of some $C(K)$-space ($K$ compact and Hausdorff). So, we can transfer problems regarding tensor products of Archimedean vector lattices with order unit in terms of $C(K)$-spaces. However, the issue is that Archimedean vector lattices with order units are scarce in the category of all vector lattices. On the other hand, by the Maeda-Ogasawara theorem ([1, Theorem 7.29]), every Archimedean vector lattice can be considered as an order dense vector sublattice of some $C^{\infty}(X)$-space for some compact Hausdorff extremally disconnected topological space $X$. Now, if we want to develop a similar theory for the tensor product between arbitrary Archimedean vector lattices in terms of $C^{\infty}(X)$-spaces, the issue is that the Cartesian product of two extremally disconnected topological spaces $X$ and $Y$, is not extremally disconnected, in general. So, $C^{\infty}(X\times Y)$ may not a vector lattice (see [13, Example 3.8]). This problem was remarkably solved by Buskes and Wickstead in 2017 ([6]) by introducing a new larger vector lattice.
Suppose $\Sigma$ is a topological space. By $S(\Sigma)$, we mean the space of all equivalence classes of continuous functions defined on open dense subsets of $\Sigma$ under the equivalence relation $f\thicksim g$ if they coincide on the intersection of their domains. This space is introduced by Buskes and Wickstead in [6]. It is a vector lattice under the pointwise lattice and vector operations.
Note that elements of a $C^{\infty}(X)$-space ($X$ a topological space) are functions, nevertheless, while we are working with a $S(\Sigma)$-space ($\Sigma$ a topological space), we are dealing with some classes of functions in which the class of each function, contains all of functions that are equal on the intersection of their domains. This defines an equivalence relation on $S(\Sigma)$. The proof of the following lemma follows from the equivalence relation defined on $S(\Sigma)$.
Lemma 1.
Suppose $\Sigma$ is a topological space and $a\in S(\Sigma)$. If $f,g\in a$ and $t\in Dom(f)\cap Dom(g)$ for some $t\in\Sigma$, then $f(t)=g(t)$.
Now, we consider a convention: for each $a\in S(\Sigma)$, put ${Dom}(a):=\cup_{f\in a}{\text{D}om}(f)$; this is well-defined. By $a(t)=m$ also we mean there exists $f\in a$ such that $f(t)=m$. By [18, Lemma 2.2], we see that for each $f\in S(\Sigma)$, there exists an element in its class with the maximal domain so that when picking a representation for an equivalence class in $S(\Sigma)$, we may assume that it is the one with maximal domain. Therefore, for a continuous real-valued function $f$ defined on an open dense subset $\Omega_{f}$ of $\Sigma$, we can identify $f$ with its equivalence class $[f]$. With this convention, assume that $X$ is an extremally disconnected topological space. Then by restriction of every $f\in C^{\infty}(X)$ to the preimage of the reals which is an open dense subset of $X$, ([1, Definition 7.26]), we see that $C^{\infty}(\Sigma)\subseteq S(\Sigma)$. Moreover, every continuous function $f\in C(\Sigma)$ can be identified with $[f]\in S(\Sigma)$, in which, $[f]$ is the class consisting of all $g\in S(\Sigma)$ with $f\thicksim g$. Therefore, we can have the inclusion $C(\Sigma)\subseteq S(\Sigma)$.
When $\Sigma$ is extremally disconnected, there is a nice relation between $S(\Sigma)$ and $C^{\infty}(\Sigma)$.
Lemma 2.
Suppose $\Sigma$ is an extremally disconnected topological space. Then there exists a lattice isomorphism between $S(\Sigma)$ and $C^{\infty}(\Sigma)$.
Proof.
By [1, Lemma 7.25], for each $f\in S(\Sigma)$, there exists a unique extension $f^{*}\in C^{\infty}(\Sigma)$. Define $\Phi:S(\Sigma)\to C^{\infty}(\Sigma)$ defined via $\Phi(f)=f^{*}$. We show that $\Phi$ is a lattice isomorphism. First, note that $\Phi$ is well-defined as the extension is unique. $\Phi$ is linear. Suppose $f:\Omega_{1}\to\mathbb{R}$ and $g:\Omega_{2}\to\mathbb{R}$ are elements of $S(\Sigma)$. Put $\Omega=\Omega_{1}\cap\Omega_{2}$ which is an open dense subset of $\Sigma$, again. Note that $f^{*}+g^{*}$ is an extension of $f+g$ defined on $\Omega$. So, by uniqueness of the extension, we have $\Phi(f+g)=(f+g)^{*}=f^{*}+g^{*}=\Phi(f)+\Phi(g)$. Now, assume that $h\in C^{\infty}(\Sigma)$. By the definition of the extended functions, $\Omega_{h}=h^{-1}(\mathbb{R})=\{x\in\Sigma,-\infty<h(x)<\infty\}$ is an open dense subset of $\Sigma$ so that the restriction of $h$ to $\Omega_{h}$ (denoted by $h_{0}$) lies in $S(\Sigma)$. It is easy to see that $\Phi(h_{0})=h$. Therefore, $\Phi$ is onto. Suppose, $\Phi(f)=f^{*}=0$ so that for each $t\in\Omega_{1}$, we have $f(t)=0$. Therefore, $\Phi$ is also one-to-one. Moreover, for each $t\in\Omega$, we have, $(f\vee g)^{*}(t)=(f\vee g)(t)=f(t)\vee g(t)=f^{*}(t)\vee g^{*}(t)$. So, again, uniqueness of the extension, results in $\Phi(f\vee g)=(f\vee g)^{*}=f^{*}\vee g^{*}=\Phi(f)\vee\Phi(g)$. This completes the proof. ∎
Considering Lemma 2 with [1, Theorem 7.27], we have the following.
Corollary 3.
Suppose $\Sigma$ is an extremally disconnected topological space. Then, $S(\Sigma)$ is a universally complete vector lattice.
We have seen that $C(\Sigma)$, the space of all real-valued continuous functions on topological space $\Sigma$, can be considered as a vector sublattice of $S(\Sigma)$. But, we have more if we consider completely regular topological spaces.
Lemma 4.
Suppose $\Sigma$ is a completely regular topological space. Then, $C(\Sigma)$ is order dense in $S(\Sigma)$.
Proof.
Suppose $0\neq g\in{S(\Sigma)}_{+}$. So, there exists some $t\in Dom(g)$ such that $g(t)=r>0$. Since $g$ is continuous, there exists a non-empty open set $U\subseteq\Sigma$ containing $t$ such that $g(s)>\frac{r}{2}$ for each $s\in U\cap Dom(g)$. Since $\Sigma$ is completely regular, there is $0\neq f\in{C(\Sigma)}_{+}$ such that $f\leq r\textbf{1}$ and $f\equiv 0$ outside of $U$. So, it is easily seen that $0<f\leq g$. ∎
Now, we characterize $uo$-convergence in $S(\Sigma)$. The proof is similar to the proof of [4, Theorem 7.1]; see also [5, Theorem 3.2].
Lemma 5.
Suppose $\Sigma$ is a completely regular topological space. For a net $(f_{\alpha})\subseteq{S(\Sigma)}_{+}$, $f_{\alpha}\xrightarrow{uo}0$ if and only if for each non-empty open set $U\subseteq\Sigma$ and for each $\varepsilon>0$, there exist a non-empty open set $V\subseteq U$ and an index $\alpha_{0}$ such that $f_{\alpha}(t)\leq\varepsilon$ for each $t\in V\cap Dom(f_{\alpha})$.
Proof.
Suppose $f_{\alpha}\xrightarrow{uo}0$. Assume that $U\subseteq\Sigma$ is a non-empty open set and $\varepsilon>0$ is arbitrary. Since $\Sigma$ is completely regular, there exists a non-zero positive continuous function $g$ on $\Sigma$ with $g\leq\textbf{1}_{U}$. By [4, Theorem 6.2], we can find non-zero positive function $h\in S(\Sigma)$ and an $\alpha_{0}$ such that $(f_{\alpha}-\varepsilon h)^{+}\perp g$ for each $\alpha\geq\alpha_{0}$. Therefore, $(f_{\alpha}-\varepsilon h)^{+}$ vanishes on non-empty open subset $V:=supp(g)\subseteq U$. Thus, for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$, $f_{\alpha}(x)\leq\varepsilon$.
For the other direction, Take any non-zero positive $h\in S(\Sigma)$. There are non-empty open set $U\subseteq\Sigma$ and $\varepsilon>0$ with $h\geq\varepsilon\textbf{1}_{U}$. By the assumption, There are non-empty open set $V\subseteq U$ and an $\alpha_{0}$ such that $f_{\alpha}(x)\leq\varepsilon$ for each $x\in V\cap Dom(f_{\alpha})$ and for each $\alpha\geq\alpha_{0}$. Since $\Sigma$ is completely regular, there is a non-zero positive continuous function $g$ with $g\leq\varepsilon\textbf{1}_{U}$. Now, we can see that $(f_{\alpha}-h)^{+}$ vanishes on $V$ so that disjoint with $g$. Again, using [4, Theorem 6.2], convinces us that $f_{\alpha}\xrightarrow{uo}0$. ∎
Moreover, we have the following standard facts. We present the proof for the sake of completeness.
Lemma 6.
Suppose $\Sigma$ and $\Omega$ are completely regular topological spaces. Then, $S(\Sigma)\overline{\otimes}S(\Omega)$ is order dense in $S(\Sigma\times\Omega)$.
Proof.
Suppose $h\in{S(\Sigma\times\Omega)}_{+}$. There exists $(t_{0},s_{0})\in\Sigma\times\Omega$ such that $h(t_{0},s_{0})=r>0$. Since $h$ is continuous, there exist non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $h(t,s)>\frac{r}{2}$ for all $(t,s)\in U\times V$. Since $\Sigma$ and $\Omega$ are completely regular, we can find $0\neq f\in{S(\Sigma)}_{+}$ such that $f\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Sigma)}$ and $f$ vanishes outside of $U$. Also, there is $0\neq g\in{S(\Omega)}_{+}$ with $g\leq\sqrt{\frac{r}{2}}{\mathbf{1}}_{S(\Omega)}$ and $g=0$ outside of $V$. We show that $0<f\otimes g\leq h$. Since both $f$ and $g$ are non-zero, $f\otimes g>0$. Suppose $(t,s)\in(Dom(f)\times Dom(g))\cap Dom(h)$. First assume that $(t,s)\in U\times V$. Then, $(f\otimes g)(t,s)=f(t)g(s)\leq\frac{r}{2}<h(t,s)$. If $(t,s)\notin U\times V$, then either $f(t)=0$ or $g(s)=0$. Therefore, $(f\otimes g)(t,s)=f(t)g(s)=0\leq h(t,s)$. ∎
The proof of the following result follows from [13, Proposition 3.2].
Lemma 7.
Suppose $\Sigma$ and $\Omega$ are topological spaces. Furthermore, assume that $E$ is an order dense vector sublattice of $S(\Sigma)$ and $F$ is an order dense vector sublattice of $S(\Omega)$. Then, $E\overline{\otimes}F$ is order dense in $S(\Sigma)\overline{\otimes}S(\Omega)$.
Theorem 8.
Suppose $E$ and $F$ are Archimedean vector lattices. Moreover, assume that $f_{\alpha}\xrightarrow{uo}f$ in $E$ and $g_{\alpha}\xrightarrow{uo}g$ in $F$. Then, $f_{\alpha}\otimes g_{\alpha}\rightarrow f\otimes g$ in the Fremlin tensor product $E\overline{\otimes}F$.
Proof.
First, assume that $(f_{\alpha}),(g_{\alpha})\geq 0$. By considering [12, Lemma 3.1], we conclude that $f,g\geq 0$. Note that by the known Maeda-Ogasawara theorem ([1, Theorem 7.29]) and using Lemma 2, There are two compact Hausdorff extremally disconnected topological spaces $\Sigma$ and $\Omega$ such that $E$ is an order dense vector sublattice of $S(\Sigma)$ and $F$ is an order dense vector sublattice of $S(\Omega)$. Thus, by [1, Theorem 1.23], $E$ and $F$ are regular sublattices in $S(\Sigma)$ and $S(\Omega)$, respectively. So, by [11, Theorem 3.2], $f_{\alpha}\xrightarrow{uo}f$ in $S(\Sigma)$ and $g_{\alpha}\xrightarrow{uo}g$ in $S(\Omega)$. On the other hand, by [6, Proposition 3.1], $E\overline{\otimes}F$ can be considered as a vector sublattice of $S(\Sigma\times\Omega)$ generated by the mappings $(h\otimes k)(t,s)=h(t)k(s)$ for each $h\in S(\Sigma)$ and for each $k\in S(\Omega)$. Compatible with [5, Remark 4.1] and Lemma 2, we show that $f_{\alpha}\otimes g_{\alpha}\xrightarrow{uo}f\otimes g$ in $S(\Sigma\times\Omega)$. Note that the constant one function $\textbf{1}_{S(\Sigma)}$ is a weak unit for $S(\Sigma)$ and $\textbf{1}_{S(\Omega)}$ is a weak unit for $S(\Omega)$ so that $\textbf{1}_{S(\Sigma)}\otimes\textbf{1}_{S(\Omega)}$ is a weak unit for $S(\Sigma)\overline{\otimes}S(\Omega)$ by [3, Proposition 4.4] so that in $S(\Sigma\times\Omega)$ by Lemma 6. So, it is enough to consider $uo$-convergence just in terms of weak units by [11, Corollary 3.5]. Now, we use Lemma 5. Suppose $W\subseteq\Sigma\times\Omega$ is a non-empty open set and $\varepsilon\in(0,1)$ is arbitrary. There are some non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $U\times V\subseteq W$. We can find non-empty open sets $U_{1}\subseteq U$ and $V_{1}\subseteq V$ with $(f_{\alpha}-f)(t)<\varepsilon$ for each $t\in U_{1}\cap Dom(f_{\alpha})\cap Dom(f)$ and $(g_{\alpha}-g)(s)<\varepsilon$ for each $s\in V_{1}\cap Dom(g_{\alpha})\cap Dom(g)$, provided that $\alpha$ is sufficiently large. Pick any $t_{0}$ in non-empty open set $U_{1}\cap Dom(f)$. Since $f$ is continuous at $t_{0}$ (WOLG, $f(t_{0})=1)$), there is a non-empty open set $U_{2}\subseteq U_{1}$ (containing $t_{0}$) such that $f(t)\leq 1$ for each $t\in U_{2}\cap Dom(f)$. Similarly, there exists a non-empty open set $V_{2}\subseteq V_{1}$ such that $g(s)\leq 1$ for each $s\in V_{2}\cap Dom(g)$. Then for each $(t,s)\in(U_{2}\times V_{2})\cap((Dom(f_{\alpha})\cap Dom(f))\times(Dom(g_{\alpha})\cap Dom(g)))=(U_{2}\cap Dom(f_{\alpha})\cap Dom(f))\times(V_{2}\cap Dom(g_{\alpha})\cap Dom(g)))$ and for sufficiently large $\alpha$, we have
$(f_{\alpha}\otimes g_{\alpha}-f\otimes g)(t,s)=(f_{\alpha}\otimes(g_{\alpha}-g))(t,s)+((f_{\alpha}-f)\otimes g)(t,s)$
$=f_{\alpha}(t)(g_{\alpha}-g)(s)+(f_{\alpha}-f)(t)g(s)<(\varepsilon+2)\varepsilon.$
Now, an easy application of Lemma 6 and Lemma 7 convinces us that $f_{\alpha}\otimes g_{\alpha}\xrightarrow{uo}f\otimes g$ in $E\overline{\otimes}F$.
For the general case, observe that we can write $f_{\alpha}={f_{\alpha}}^{+}-{f_{\alpha}}^{-}$ and $g_{\alpha}={g_{\alpha}}^{+}-{g_{\alpha}}^{-}$. Moreover, by using [12, Lemma 3.1], we see that $f_{\alpha}\xrightarrow{uo}f$ if and only if ${f_{\alpha}}^{\pm}\xrightarrow{uo}{f}^{\pm}$ and also $g_{\alpha}\xrightarrow{uo}g$ if and only if ${g_{\alpha}}^{\pm}\xrightarrow{uo}{g}^{\pm}$. By using the former case, we have the following observation.
$f_{\alpha}\otimes g_{\alpha}=({f_{\alpha}}^{+}-{f_{\alpha}}^{-})\otimes({g_{\alpha}}^{+}-{g_{\alpha}}^{-})=({f_{\alpha}}^{+}\otimes{g_{\alpha}}^{+})-({f_{\alpha}}^{+}\otimes{g_{\alpha}}^{-})-({f_{\alpha}}^{-}\otimes{g_{\alpha}}^{+})+$
$({f_{\alpha}}^{-}\otimes{g_{\alpha}}^{-})\xrightarrow{uo}(f^{+}\otimes g^{+})-(f^{+}\otimes g^{-})-(f^{-}\otimes g^{+})+(f^{-}\otimes g^{-})=f\otimes g.$
Proposition 9.
Suppose $E$ and $F$ are vector lattices. Assume that $(x_{\alpha})\subseteq E$ is $uo$-null and $(y_{\alpha})\subseteq F$ is eventually order bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uo}0$ in $E\overline{\otimes}F$.
Proof.
There exists $y\in F_{+}$ such that $|y_{\alpha}|\leq y$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Now, we use the following inequality.
$|x_{\alpha}\otimes y_{\alpha}|\wedge w\leq(|x_{\alpha}|\otimes|y_{\alpha}|)\wedge(x_{0}\otimes y_{0})$
$\leq(|x_{\alpha}|\wedge x_{0})\otimes(|y_{\alpha}|\vee y_{0})\leq(|x_{\alpha}|\wedge x_{0})\otimes(y\vee y_{0}).$
Now, by [13, Corollary 3.4], we conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uo}0$, as claimed. ∎
Remark 10.
Note that order bounded assumption is essential in Proposition 9 and can not be removed. It is know that in $L_{p}(\mu)$-spaces, for sequences, $uo$-convergence and almost every where convergence agree. Put $E=L^{2}(\mathbb{R})$, $f_{n}(x)=\frac{x}{n}$ and $g_{n}(x)=n$. It is easy to see that $(f_{n})$ is $uo$-null and $(g_{n})$ is not order bounded. Moreover, $(f_{n}\otimes g_{n})_{n}$ is the identity mapping which is not pointwise convergent to zero, certainly. More precisely, by [10, Proposition 4B], $L^{2}(X)\widehat{\otimes}L^{2}(Y)$ can be considered as vector sublattice of $L^{2}(X\times Y)$, where $X$ and $Y$ are measure spaces. Moreover, $L^{2}(X)\overline{\otimes}L^{2}(Y)$ is an order dense in $L^{2}(X)\widehat{\otimes}L^{2}(Y)$.
It can be easily verified that if a net $(x_{\alpha})\subseteq E$ and a net $(y_{\alpha})\subseteq F$ are eventually order bounded, then, the net $(x_{\alpha}\otimes y_{\alpha})$ is also eventually order bounded in $E\overline{\otimes}F$. So, we have the following fact that is an extension of [13, Corollary 3.4], as well.
Corollary 11.
[…]
3.2. Fremlin projective tensor product and unbounded convergences
[…]
Lemma 12.
Suppose $E$ and $F$ are Banach lattices such that the Fremlin projective tensor product $E\widehat{\otimes}F$ is order continuous. Then, both $E$ and $F$ are order continuous.
Proof.
We show that $E$ is order continuous; order continuity of $F$ is similar. Suppose on a contrary, $E$ is not order continuous. So, there exists a disjoint order bounded sequence $(x_{n})\subseteq E$ which is not norm null. Choose any non-zero positive $y\in F$. It is easy to see that the sequence $(x_{n}\otimes y)$ is disjoint and order bounded in $E\widehat{\otimes}F$. By the assumption, it is norm null. But, the projective tensor norm is a cross-norm so that $(x_{n})$ should be norm null that is a contradiction. ∎
Observe that the converse of Lemma 12 is not true even for very nice spaces; as shown in [10, example 4C]. More precisely, put $E=F=L^{2}[0,1]$. Then, the Fremlin projective tensor product $E\widehat{\otimes}F$ is not order complete so that the projective norm on $E\widehat{\otimes}F$ is not order continuous, although both $E$ and $F$ are order continuous (even reflexive) Banach lattices.
Proposition 13.
Suppose $E$ and $F$ are Banach lattices. Assume that $(x_{\alpha})\subseteq E$ is $un$-null and $(y_{\alpha})\subseteq F$ is eventually norm bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}0$ in $E\widehat{\otimes}F$.
Proof.
By the assumption, $\|y_{\alpha}\|\leq 1$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Now, we use the following inequality.
$|x_{\alpha}\otimes y_{\alpha}|\wedge w\leq|x_{\alpha}|\otimes|y_{\alpha}|\wedge x_{0}\otimes y_{0}$
$\leq(|x_{\alpha}|\wedge x_{0})\otimes(|y_{\alpha}|\vee y_{0}).$
Therefore,
$\||x_{\alpha}\otimes y_{\alpha}|\wedge w\|\leq\|\leq\||x_{\alpha}|\wedge x_{0}\|\||y_{\alpha}|\vee y_{0}\|\rightarrow 0.$
We conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}0$ in $E\overline{\otimes}F$ and so that in $E\widehat{\otimes}F$ by [14, Theorem 4.3]; see also [16, Lemma 3.5]. ∎
Remark 14.
Note that norm boundedness is essential in Proposition 13 and can not be omitted. Put $E=\ell_{\infty}$. Consider two sequences $(u_{n})$ and $(v_{n})$ in $E$ defined via $u_{n}=(0,\ldots,0,n,0,\ldots)$ and $v_{n}=(0,\dots,0,\frac{1}{n},0,\ldots)$. Indeed, $v_{n}\xrightarrow{un}0$ and $(u_{n})$ is unbounded. Consider the Fremlin tensor product $E\widehat{\otimes}E$. The sequence $(u_{n}\otimes v_{n})_{n\in\mathbb{N}}$ is the sequence $(e_{n})$ consisting of the standard basis of $E$; that is not $un$-null by [14, Theorem 3.2] (note that $E\widehat{\otimes}E$ is a Banach lattice with order unit).
Theorem 15.
Suppose $E$ and $F$ are Banach lattices such that $E\widehat{\otimes}F$ is order continuous. If $(x_{\alpha})\subseteq E$ is $un$-convergent to $x\in E$ and $(y_{\alpha})\subseteq F$ is $un$-convergent to $y\in F$, then $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}x\otimes y$ in $E\widehat{\otimes}F$.
Proof.
First note that by Lemma 12, both $E$ ad $F$ are order continuous. Moreover, observe that by [8, Corollary 3.5], there exists an increasing sequence $(\alpha_{k}^{\prime})$ of indices such that $x_{\alpha_{k}^{\prime}}-x\xrightarrow{un}0$. By [8, Theorem 3.2], there exists a subsequence $(\alpha_{k})$ of $(\alpha_{k}^{\prime})$ and a disjoint sequence $(d_{k})$ in $E$ such that $\|x_{\alpha_{k}}-x-d_{k}\|\rightarrow 0$. By Proposition 13, $(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)\xrightarrow{un}0$, $(x_{\alpha_{k}}-x)\otimes y\xrightarrow{un}0$ and $x\otimes(y_{\alpha}-y)\xrightarrow{un}0$. Furthermore, observe that the net $d_{k}\otimes(y_{\alpha}-y)$ is disjoint in $E\widehat{\otimes}F$ so that $un$-null by [14, proposition 3.5]. Therefore, we have
$x_{\alpha_{k}}\otimes y_{\alpha}-x\otimes y=(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)+d_{k}\otimes(y_{\alpha}-y)+$
$x\otimes(y_{\alpha}-y)+(x_{\alpha_{k}}-x)\otimes y\xrightarrow{un}0.$
This procedure happens for every subnet of $(x_{\alpha})$ and $(y_{\alpha})$; since $un$-convergence is topological, we conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}x\otimes y$. ∎
Question 3.
Does the result of Theorem 15 hold without order continuity of the Fremlin projective tensor product $E\widehat{\otimes}F$.
Now, we investigate the relation between unbounded absolute weak convergence and the Fremlin projective tensor product.
Proposition 16.
Suppose $E$ and $F$ are Banach lattices. Assume that $(x_{\alpha})\subseteq E$ is $uaw$-null and $(y_{\alpha})\subseteq F$ is eventually order bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}0$ in $E\widehat{\otimes}F$.
Proof.
By the assumption, there exists $y_{1}\in F_{+}$ such that $y_{\alpha}\leq y_{1}$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Fix $f\in(E\widehat{\otimes}F)^{{}^{\prime}}_{+}=B^{r}(E\times F)_{+}$, where $B^{r}(E\times F)$ is the Banach lattice of all bounded bilinear forms defined on $E\times F$; see [10, 1I] for more details. Now, we use the following inequality.
$|f(x_{\alpha}\otimes y_{\alpha}\wedge w)|\leq f(|x_{\alpha}|\otimes|y_{\alpha}|\wedge x_{0}\otimes y_{0})$
$\leq f(|x_{\alpha}|\wedge x_{0},|y_{\alpha}|\vee y_{0})\leq f(|x_{\alpha}|\wedge x_{0},y_{0}\vee y_{1}).$
Note that the restriction $f_{y_{0}\vee y_{1}}:E\to\mathbb{R}$ is a positive linear functional. Therefore, by the assumption,
$f(|x_{\alpha}\otimes y_{\alpha}|\wedge w)\leq f(|x_{\alpha}\wedge x_{0},y_{0}\vee y_{1})\rightarrow 0.$
We conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}0$ in $E\overline{\otimes}F$ and so that in $E\widehat{\otimes}F$ by [16, Lemma 3.4]. ∎
Theorem 17.
Suppose $E$ and $F$ are Banach lattices such that of one them is order continuous. If $(x_{\alpha})\subseteq E$ is $uaw$-convergent to $x\in E$ and $(y_{\alpha})\subseteq F$ is $uaw$-convergent to $y\in F$, then $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}x\otimes y$ in $E\widehat{\otimes}F$.
Proof.
Assume that $E$ is order continuous. By [19, Theorem 4], $x_{\alpha}\xrightarrow{un}x$. By [8, Theorem 3.2], there exists a increasing sequence $(\alpha_{k})$ of indices and a disjoint sequence $(d_{k})$ in $E$ such that $\|x_{\alpha_{k}}-x-d_{k}\|\rightarrow 0$. By passing to a further subsequence, we may assume that $(x_{\alpha_{k}}-x-d_{k})$ is order bounded. So, by Proposition 16, $(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)\xrightarrow{uaw}0$, $(x_{\alpha_{k}}-x)\otimes y\xrightarrow{uaw}0$ and $x\otimes(y_{\alpha}-y)\xrightarrow{uaw}0$. Note that the net $(d_{k}\otimes(y_{\alpha}-y))$ is disjoint in $E\widehat{\otimes}F$ so that $uaw$-null by [19, Lemma 2]. Thus, we conclude that
$x_{\alpha_{k}}\otimes y_{\alpha}-x\otimes y=(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)+d_{k}\otimes(y_{\alpha}-y)+$
$x\otimes(y_{\alpha}-y)+(x_{\alpha_{k}}-x)\otimes y\xrightarrow{uaw}0.$
This argument happens for every subnet; therefore, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}x\otimes y$. ∎
Question 4.
Is order continuity of one of the components necessary as an assumption for Theorem 17?
Acknowledgments. Most of the work on this paper was done during a visit of the author to University of Alberta in 2023. This note would not have existed without invaluable help and discussions of V. G. Troitsky, my friend and my colleague. The author would like to thank him so much for great hospitality during his visit, as well. Thanks is also due to Eugene Bilokopytov for many useful discussions.